OpenStudy (anonymous):

rationalize the denominator of 5 / (6 + (3 sqrt(2)) - (2 + (2 sqrt(2))

OpenStudy (lalaly):

\[\frac{5}{6+3\sqrt{2}-2+2\sqrt2}\] \[=\frac{5}{4+5\sqrt2} \times \frac{4-5\sqrt2}{4-5\sqrt2}\]\[=\frac{20-25\sqrt2}{6}\]